In 1995, Sweden and Norway had more women members of parliament (MPs) than any other democratic countries. Combined, the two countries had 206 women in parliaments. Taking double the number of women MPs in Sweden and adding this to triple the number of women MPs in Norway results in 477.
a. How many women MPs are there in each country?
b. The total MPs in Sweden is 349 and in Norway the number is 165. What percent (to the nearest whole percent) of the total MPs are women for each of the two countries?

Answer by Theo(13342)

let s = number of women members of parliament in sweden.
let n = number of women members of parliament in norway.

s + n = 206

2s + 3n = 477

solve for s in the first equation to get s = 206 - n

substitute in the second equation to get 2*(206-n) + 3n = 477

simplify to get 412 - 2n + 3n = 477

combine like terms to get 412 + n = 477

subtract 412 from both sides of equation to get n = 477 - 412 = 65

if n = 65, then s = 206 - 65 = 141

equation of s + n = 206 becomes 141 + 65 = 206 which is true.

equation of 2s + 3n = 477 becomes 2*141 + 3*65 = 477 becomes 282 = 195 = 477 which is also true.

numbers look good.

s = 141 = number of women members of parliament in sweden.
n = 65 = number of women members of parliament in norway.

total members of parliament in sweden is 349.
total women members of parliament in sweden is 141.
141/349 = .404011461 = 40% to the nearest whole percent.

total members of parliament in norway is 165.
total women members of parliament in norway is 65.
65/165 = .393939394 = 39% to the nearest whole percent.